OpenStudy (anonymous):

Can anyone just give me a list of fomulas? And maybe what they are used for?

OpenStudy (anonymous):

Translational motion x = x 0 + v 0 t + 1/2at2 Vƒ = Vo + at Frictional force f max = μ Ν μk < μs always Uniform circular motion Fc = mac = mv2 /r ac= v2 /r Momentum, Impulse I = F Δt = ΔM M = mv Work, Power W = F d cosθ P = ΔW/Δt Energy (conservation) ET = Ek + Ep E = mc2 Spring Force, Work F = -kx W = kx2 /2 Continuity (fluids) A v = const. ρAv = const. Current and Resistance I = Q/t R = ρl/A Resistors (series, par.) Req = R1 + R2 . . . 1/ Req = 1/ R1 +1/ R2 Sound dB = 10 log 10 (I/I0 ) beats = Δ f Kirchoff's Laws Σi = 0 at a junction ΣΔV = 0 in a loop Thermodynamics Q = mc Δ T (MCAT !) Q = mL Torque forces L1 = F1× r1 (CCW + ve) L2 = F2 × r2 (CW -ve) Torque force at EQ ΣFx = 0 and ΣFy = 0 ΣL = 0 Refraction ( sin θ1 )/(sin θ2 ) = v1 /v2 = n2 /n1 = λ1 /λ2 n = c/v

OpenStudy (anonymous):

Is there anything.... Simpler?...

OpenStudy (anonymous):

That's too much formula.

OpenStudy (anonymous):

Well, you asked for a list of formulae.... What were you expecting? Is there a particular list of formulae that you are looking for, I mean that includes most of the formulae for basic physics.